Astro:

As part of a school science project, Caroline needs to make a model of planet Earth with a 7-inch diameter using playdough. If one container of playdough contains 9 cubic inches, how many containers will she need?

YECR:

What do you think you need to do? Divide Or Multipy the 7 and 9?

Astro:

I'm not sure.

snowflake0531:

We first have to find volume of the sphere \[\frac {4} {3} \pi r^3\]

snowflake0531:

So diameter is 2 times radius so the radius in this question is...?

Astro:

3.5?

snowflake0531:

Yep so we have \[\frac {4} {3} \pi ~(3.5^3)\] can you find the decimal form of that?

Astro:

178.76?

snowflake0531:

I got 179.59 approximately-?

1 attachment
Astro:

Oh- Okay I see now.

snowflake0531:

So then divide that by 9, and round to the next whole number, even if it isn't supposed to round up, theoretically

Astro:

If its rounded up would it be 20?

snowflake0531:

yep, so your answer is 20

Astro:

Thank you again (:

snowflake0531:

yw~ xd